OpenStudy (anonymous):

(6n-1)^2+2(6n-1)-80=0 n=??

OpenStudy (anonymous):

Lets do variable substitution. Let x = 6n-1. Then your equation becomes: \[x^{2} +2x-80 = 0\] This factors nicely: \[(x-8)(x+10) = 0\] So x = 8 and x = -10 Now to put back out substitution: 6n -1 = 8 means n = 3/2 6n-1 = -10 means n = -3/2

OpenStudy (anonymous):

wow! thank you! we haven't even learned about variable substitution but i like it!